show the value
f(x)=-3(x+2)^2+3

Answer by Nate(3500)

f(x)=-3(x+2)^2+3
y=-3(x+2)^2+3
this is vertex form: y=a(x-h)^2+k
the vertex is: (h,k) or (-2,3)
when the value of x is -2 and the y value is 3